OpenStudy (anonymous):

Use basic identities to simplify the expression. cosecant of theta times cotangent of theta divided by secant of theta.

OpenStudy (anonymous):

write everything in terms of sine and cosine and see what cancels

OpenStudy (anonymous):

\( \huge \frac{\csc \theta \times \cot \theta}{\sec \theta} \) This is what we are working with correct?

OpenStudy (anonymous):

These things can get tricky but they are fun, when you are not being timed to do them lol

OpenStudy (anonymous):

Haha yeah and that is what the question is :)

OpenStudy (decentnabeel):

@Nixy cot=cos/sin not sin/cos

OpenStudy (anonymous):

opps yep thanks for the catch

OpenStudy (anonymous):

\( \huge \frac{\frac{1}{\sin \theta} \times \frac{cos}{sin}}{\sec \theta} \) now work the top

OpenStudy (anonymous):

Uhhh this is bad, I don't know how to....sorry I am taking this online and my teacher isn't that helpful :(

OpenStudy (decentnabeel):

\[\mathrm{Simplify}\:\frac{\frac{1}{\sin \left(θ\right)}\cdot \frac{\sin \left(θ\right)}{\cos \left(θ\right)}}{\sec \left(θ\right)}:\quad 1\]

OpenStudy (anonymous):

wait isn't it cos over sin?

OpenStudy (anonymous):

Yes, he made a boo boo too :-D

OpenStudy (decentnabeel):

yes

OpenStudy (anonymous):

sec is 1/cos theta correct? so would I put that in?

OpenStudy (decentnabeel):

yes

OpenStudy (anonymous):

and then cancel out the cos?

OpenStudy (anonymous):

@peachpi chime in any time to help :-) @peachpi is good at these. She/He has helped me on some before.

OpenStudy (anonymous):

Oh okay thanks :)

OpenStudy (anonymous):

you guys look like you have it under control

OpenStudy (anonymous):

write the denominator as 1/cos Θ. then flip it to the reciprocal when you change division to multiplication